# EOV-Seg: Efficient Open-Vocabulary Panoptic Segmentation

### v1

@@ Abstract
Open-vocabulary panoptic segmentation aims to segment and classify everything in diverse scenes across an unbounded vocabulary. Existing methods typically employ two-stage or single-stage framework. The two-stage framework involves cropping the image multiple times using masks generated by a mask generator, followed by feature extraction, while the single-stage framework relies on a heavyweight mask decoder to make up for the lack of spatial position information through self-attention and cross-attention in multiple stacked Transformer blocks. Both methods incur substantial computational overhead, thereby hindering the efficiency of model inference. To fill the gap in efficiency, we propose EOV-Seg, a novel single-stage, shared, efficient, and spatial-aware framework designed for open-vocabulary panoptic segmentation. Specifically, EOV-Seg innovates in two aspects. First, a Vocabulary-Aware Selection (VAS) module is proposed to improve the semantic comprehension of visual aggregated features and alleviate the feature interaction burden on the mask decoder. Second, we introduce a Two-way Dynamic Embedding Experts (TDEE), which efficiently utilizes the spatial awareness capabilities of ViT-based CLIP backbone. To the best of our knowledge, EOV-Seg is the first open-vocabulary panoptic segmentation framework towards efficiency, which runs faster and achieves competitive performance compared with state-of-the-art methods. Specifically, with COCO training only, EOV-Seg achieves 24.2 PQ, 31.6 mIoU, and 12.7 FPS on the ADE20K dataset for panoptic and semantic segmentation tasks and the inference time of EOV-Seg is 4-21 times faster than state-of-the-art methods. Especially, equipped with ResNet-50 backbone, EOV-Seg runs 25 FPS with only 71M parameters on a single RTX 3090 GPU. Code is available at https://github.com/nhw649/EOV-Seg.
@@ Introduction
Extensive experiments show that EOV-Seg runs faster and achieves competitive performance compared with state-of-the-art methods. In particular, with COCO training only, EOV-Seg achieves 24.2 PQ, 31.6 mIoU, and 12.7 FPS on the ADE20K dataset. When taking ResNet-50 as backbone, it runs 25 FPS with only 71M parameters on a single RTX 3090 GPU.
@@
Architecture. We use a CNN-based CLIP (Radford et al. 2021) visual encoder as backbone, S, M, L denote ResNet-50, ResNet-50x4, ConvNeXt-L respectively. Additionally, we utilize the first block of the ViT-B/16 CLIP model as a spatial awareness extractor.
Training Strategy. Following prior works  (Xu et al. 2023a; Yu et al. 2023), the training batch size is $16$. We train our EOV-Seg on $4$ NVIDIA 3090 GPUs for a total of $20k$ iterations. We only use the COCO Panoptic (Lin et al. 2014) dataset for training, with a crop size of $1024\times 1024$. All of our experiments are performed three times and averaged for fair comparison.
@@
Open-Vocabulary Panoptic Segmentation. Tab. 1 reports the performance of different open vocabulary panoptic segmentation methods on the ADE20k dataset. Compared to existing state-of-the-art methods, EOV-Seg (L) runs faster and achieves competitive performance. Specifically, although ODISE (Xu et al. 2023a) take powerful text-to-image diffusion (Rombach et al. 2022) model as backbone and is trained on the additional COCO Caption (Chen et al. 2015) dataset, EOV-Seg (L) outperforms ODISE by $+0.8$ PQ with faster speed (21$\times$), fewer parameters (-1297M), and lower computational complexity (-601GFLOPs). Compared to FC-CLIP (Yu et al. 2023), EOV-Seg (L) is 4 times faster than it and still shows competitive performance. It is worth mentioning that compared to other state-of-the-art methods like MasQCLIP (Xu et al. 2023c), OPSNet (Chen et al. 2023), FreeSeg (Qin et al. 2023), and MaskCLIP (Zheng Ding 2023), EOV-seg shows significant improvements, such as PQ increased by $+0.7$ to $+9.1$, SQ increased by $+0.5$ to $+17.7$, RQ increased by $+0.8$ to $+10.8$, mIoU increased by $+1.3$ to $+7.9$ and the inference speed is $3.8$ to $18.1$ times faster. These results demonstrate that our method achieves the optimal trade-off between performance and speed.
Task-general EOV-Seg Is All You Need. Our method is task-universal, and it can not only perform panoptic segmentation but also semantic segmentation. In Tab. 2, we evaluate the open-vocabulary semantic segmentation performance on ADE20K (Zhou et al. 2017) with 150 classes (A-150) and 847 classes (A-847), PASCAL Context (Mottaghi et al. 2014) with 459 classes (PC-459) and 59 classes (PC-59), as well as PASCAL VOC (Everingham et al. 2010) (PAS-20) datasets. EOV-Seg delivers performance comparable to current state-of-the-art models and it boasts the fastest inference speed among its peers. Specially, EOV-Seg (L) outperforms CAT-Seg (Cho et al. 2024) by +2.1 mIoU on the A-150 dataset and is 6$\times$ faster, has fewer parameters (-265M) and lower GLOPs than CAT-Seg. These results demonstrate that even though EOV-Seg is trained only on the COCO Panoptic dataset with fewer semantic classes, it successfully adapts to semantic segmentation tasks, generalizes efficiently to unknown categories, achieves accurate segmentation, and runs faster with fewer parameters.
Ablation Study
@@
DINOv2 0 21.9 67.3 27.3 29.5 11.7 229M
5 22.4 67.7 28.8 30.1 8.5 279M
11 23.1 68.2 29.4 31.2 3.8 306M
SAM 0 22.4 68.3 27.8 30.8 12.1 232M
5 23.7 70.4 28.8 31.6 8.7 275M
11 24.5 71.2 29.5 32.2 3.2 305M
ViT-based CLIP 0 24.2 70.1 30.0 31.6 12.7 225M
5 23.8 68.6 29.4 32.0 10.3 247M
@@
Spatial Awareness Extractor. In Tab. 3, we compared the performance and efficiency of using ViT-B versions of different pre-trained models as spatial-aware extractors. Specifically, when using different blocks of the DINOv2 backbone as spatial awareness extractor, it shows poor performance. As shown in Fig. 3 (b), it contains a lot of noise and lacks fine-grained spatial information, we argue the difficulty in learning feature space patterns due to self-supervised pre-training in DINOv2, coupled with the use of unlabeled data, results in diminished instance discrimination and spatial awareness capabilities. Since SAM is pre-trained on the fine-grained SA-1B (Kirillov et al. 2023) dataset, it exhibits enhanced fine-grained spatial representation capabilities in its later blocks, progressively improving the segmentation quality. However, the increase in RQ is modest, likely constrained by its class-agnostic pre-training. One key finding is that the first ViT-based block demonstrates comparable performance to SAM, with faster inference 4×4$\times$4 × and fewer parameters (-80M). This is consistent with the observation in Fig. 3 (d). This improvement is due to the ability of early self-attention mechanism to process global information, where each patch of the image captures its respective details. As the number of blocks increases, performance is saturated and has a downward trend. To achieve the best trade-off between performance and inference speed, the first block of the ViT-based CLIP is the best choice.
@@
Our Two Innovative Components. In Tab. 4, we designed two additional variants of spatial awareness fusion, namely EAF and SDI, for more details in the appendix. All three variants of spatial awareness fusion significantly improved performance, demonstrating the importance of spatial location information for the lightweight decoder. When we introduce the VAS module for EAF, the performance tends to decline. This phenomenon is attributed to the early integration of features, which changes the spatial distribution of features and leads to the loss of key spatial information in important semantic features. However, VAS can cooperate with SDI or TDEE spatial awareness fusion module for mutual benefit and win-win results, and can bring additional performance improvements. In summary, We finally choose the TDEE and VAS modules, which work together as two innovative components and bring performance improvements of +2.1 PQ, +2.9 SQ, +2.7 RQ and +1.3 mIoU without introducing additional computational burden.
Feature Interaction Methods. In Tab. 5, we explored the performance differences between Cross-Attention (CA) and Dynamic Depthwise Attention (DDA) two attention modules and their interaction with features of different scales. When using CA with multi-scale features only yields marginal improvements over our method, with a $+0.2$ increase in PQ and $+0.3$ in mIoU, but results in a significant slowdown, approximately 4×4$\times$4 × slower in FPS. Since single scale avoids the fuzzy spatial distribution caused by spatial dimension alignment, and the pre-perception of VAS can reduce the decoder interaction burden, using DDA with single-scale can bring a performance improvement of +1.5 PQ and +1.9 mIoU over DDA with multi-scale.
@@ Conclusion
We propose EOV-Seg, a framework designed for open-vocabulary panoptic segmentation. It uses VAS and TDEE to improve instance discrimination capabilities for alleviating the interaction burden on mask decoder and endow mask recognition with spatial awareness capabilities. Extensive experiments show that our EOV-Seg is effective in terms of both accuracy and speed.
@@ Appendix A Overview
In the following supplementary materials, we elaborate on the inference details in Sec. B, provide the pseudocode of the Vocabulary-Aware Selection (VAS) module in Sec. C, and delineate the pipeline of the first two fusion methods of spatial awareness fusion, namely Early Aggregation Fusion (EAF) and Separable Dynamic Interaction (SDI) in Sec. D. In Sec. E and Sec. F, we present the additional experimental results and additional qualitative results, respectively.
@@ Appendix E Additional Experimental Results
Closed-Vocabulary and Open-Vocabulary panoptic segmentation performance. In Tab. 7, we compared EOV-Seg with SOTAs on COCO panoptic segmentation dataset to evaluate the closed-vocabulary capabilities. Specifically, with only a small increase in the number of parameters, the performance of EOV-Seg significantly surpassed FreeSeg (Qin et al. 2023), with PQ improved by +15.5, SQ improved by +2.2, RQ improved by +18.4, and mIoU improved by +18 and FPS is 16×16$\times$16 × faster. Additionally, EOV-Seg outperforms ODISE (Xu et al. 2023a) by +1.3 PQ, +7.8 mIoU, and the inference speed is $x\times$ faster with fewer parameters (-1297M). These results demonstrate that EOV-Seg not only has good generalization performance in open-vocabulary scenes, but also achieves excellent performance in closed-vocabulary scenes. In Tab. 8, we present open-vocabulary panoptic segmentation results on the Cityscapes (Cordts et al. 2016) dataset. EOV-Seg surpasses ODISE (Xu et al. 2023a), PQ is improved by +, RQ is improved by + with fewer parameters (-1297M). Even under challenging street scene datasets, EOV-Seg has strong open-vocabulary recognition and segmentation capabilities.
@@
Method PQ SQ RQ mIoU FPS Params FreeSeg (Qin et al. 2023) 31.4 78.3 38.9 42.2 0.8 219M ODISE (Xu et al. 2023a) 45.6 - - 52.4 0.7 1522M EOV-Seg (L) 46.9 80.5 57.3 60.2 12.9 225M
@@
Method Scales PQ SQ RQ mIoU FPS EOV-Seg (S) (800, 1333) 14.3 54.1 17.7 20.4 17.0 (640, 2560) 15.1 57.0 18.9 21.9 25.0 EOV-Seg (L) (800, 1333) 22.9 70.2 28.3 31.3 9.1 (640, 2560) 24.2 70.1 30.0 31.6 12.7

### v2

@@ Abstract
Open-vocabulary panoptic segmentation aims to segment and classify everything in diverse scenes across an unbounded vocabulary. Existing methods typically employ two-stage or single-stage framework. The two-stage framework involves cropping the image multiple times using masks generated by a mask generator, followed by feature extraction, while the single-stage framework relies on a heavyweight mask decoder to make up for the lack of spatial position information through self-attention and cross-attention in multiple stacked Transformer blocks. Both methods incur substantial computational overhead, thereby hindering the efficiency of model inference. To fill the gap in efficiency, we propose EOV-Seg, a novel single-stage, shared, efficient, and spatial-aware framework designed for open-vocabulary panoptic segmentation. Specifically, EOV-Seg innovates in two aspects. First, a Vocabulary-Aware Selection (VAS) module is proposed to improve the semantic comprehension of visual aggregated features and alleviate the feature interaction burden on the mask decoder. Second, we introduce a Two-way Dynamic Embedding Experts (TDEE), which efficiently utilizes the spatial awareness capabilities of ViT-based CLIP backbone. To the best of our knowledge, EOV-Seg is the first open-vocabulary panoptic segmentation framework towards efficiency, which runs faster and achieves competitive performance compared with state-of-the-art methods. Specifically, with COCO training only, EOV-Seg achieves 24.5 PQ, 32.1 mIoU, and 11.6 FPS on the ADE20K dataset and the inference time of EOV-Seg is 4-19 times faster than state-of-the-art methods. Especially, equipped with ResNet50 backbone, EOV-Seg runs 23.8 FPS with only 71M parameters on a single RTX 3090 GPU. Code is available at https://github.com/nhw649/EOV-Seg.
@@ Introduction
Extensive experiments show that EOV-Seg runs faster and achieves competitive performance compared with state-of-the-art methods. In particular, with COCO training only, EOV-Seg achieves 24.5 PQ, 32.1 mIoU, and 11.6 FPS on the ADE20K dataset. When taking ResNet50 as backbone, it runs 23.8 FPS with only 71M parameters on a single RTX 3090 GPU.
@@
Experiments
Implementation Details
Architecture. We use a CNN-based CLIP (Radford et al. 2021) visual encoder as backbone, S, M, L denote ResNet50, ResNet50x4, ConvNeXt-Large, respectively. Additionally, we utilize the first block of the ViT-B/16 CLIP model as a spatial awareness extractor.
Training Strategy. Following prior works  (Xu et al. 2023a; Yu et al. 2023), the training batch size is 16. We train our EOV-Seg on 4 NVIDIA 3090 GPUs for a total of 200k iterations. We only use the COCO Panoptic (Lin et al. 2014) dataset for training, with a crop size of 1024$\times$1024. All of our experiments are performed three times and averaged for fair comparison.
@@
Open-Vocabulary Panoptic Segmentation. Tab. 1 reports the performance of different open vocabulary panoptic segmentation methods on the ADE20k dataset. Compared to existing state-of-the-art methods, EOV-Seg (L) runs faster and achieves competitive performance. Specifically, although ODISE (Xu et al. 2023a) take powerful text-to-image diffusion (Rombach et al. 2022) model as backbone and is trained on the additional COCO Caption (Chen et al. 2015) dataset, EOV-Seg (L) outperforms ODISE by $+1.1$ PQ with faster speed (19×19$\times$19 ×), fewer parameters (-1297M), and lower computational complexity (-576GFLOPs). Compared to FC-CLIP (Yu et al. 2023), EOV-Seg (L) is about $4$ times faster than it and still shows competitive performance. It is worth mentioning that compared to other state-of-the-art methods like MasQCLIP (Xu et al. 2023c), OPSNet (Chen et al. 2023), FreeSeg (Qin et al. 2023), and MaskCLIP (Zheng Ding 2023), EOV-seg shows significant improvements, such as PQ increased by $+1.0$ to $+9.4$, SQ increased by $+0.6$ to $+17.9$, RQ increased by $+0.9$ to $+10.9$, mIoU increased by $+1.8$ to $+8.4$ and the inference speed is $3.5$ to $19.3$ times faster. These results demonstrate that our method achieves the optimal trade-off between performance and speed.
Task-general EOV-Seg Is All You Need. Our method is task-universal, and it can not only perform panoptic segmentation but also semantic segmentation. In Tab. 2, EOV-Seg delivers performance comparable to current state-of-the-art models and it boasts the fastest inference speed among its peers. Specially, EOV-Seg (L) outperforms CAT-Seg (Cho et al. 2024) by $+2.0$ mIoU on the A-150 dataset and is 5.3×5.3$\times$5.3 × faster, has fewer parameters (-265M) and lower FLOPs than CAT-Seg. These results demonstrate that even though EOV-Seg is trained only on the COCO Panoptic dataset with fewer semantic classes, it successfully adapts to semantic segmentation tasks, generalizes efficiently to unknown categories, achieves accurate segmentation, and runs faster with fewer parameters.
@@
DINOv2 0 21.9 67.3 27.3 29.5 11.7 225M
5 22.4 67.7 28.8 30.1 8.9 247M
11 23.1 68.2 29.4 31.2 5.8 260M
SAM 0 22.4 68.3 27.8 30.8 11.6 225M
5 23.7 70.4 28.8 31.6 8.7 247M
11 24.9 71.2 29.5 32.5 5.2 260M
ViT-based CLIP 0 24.5 70.2 30.1 32.1 11.6 225M
5 23.8 68.6 29.4 32.0 9.1 247M
@@
Ablation Study
Spatial Awareness Extractor. In Tab. 3, we compared the performance and efficiency of using ViT-B versions of different pre-trained models as spatial-aware extractors. Specifically, when using different blocks of the DINOv2 backbone as spatial awareness extractor, it shows poor performance. As shown in Fig. 3 (b), it contains a lot of noise and lacks fine-grained spatial information, we argue the difficulty in learning feature space patterns due to self-supervised pre-training in DINOv2, coupled with the use of unlabeled data, results in diminished instance discrimination and spatial awareness capabilities. Since SAM is pre-trained on the fine-grained SA-1B (Kirillov et al. 2023) dataset, it exhibits enhanced fine-grained spatial representation capabilities in its later blocks, progressively improving the segmentation quality. However, the increase in RQ is modest, likely constrained by its class-agnostic pre-training. One key finding is that the first ViT-based block demonstrates comparable performance to SAM, with about $\times2$ faster inference speed and fewer parameters (-35M). This is consistent with the observation in Fig. 3 (d). This improvement is due to the ability of early self-attention mechanism to process global information, where each patch of the image captures its respective details. As the number of blocks increases, performance is saturated and has a downward trend. To achieve the best trade-off between performance and inference speed, the first block of the ViT-based CLIP is the best choice.
@@
Our Two Innovative Components. In Tab. 4, we designed two additional variants of spatial awareness fusion, namely EAF and SDI, for more details in the appendix. All three variants of spatial awareness fusion significantly improved performance, demonstrating the importance of spatial location information for the lightweight decoder. When we introduce the VAS module for EAF, the performance tends to decline. This phenomenon is attributed to the early integration of features, which changes the spatial distribution of features and leads to the loss of key spatial information in important semantic features. However, VAS can cooperate with SDI or TDEE spatial awareness fusion module for mutual benefit and win-win results, and can bring additional performance improvements. In summary, We finally choose the TDEE and VAS modules, which work together as two innovative components and bring performance improvements of $+2.4$ PQ, $+3.0$ SQ, $+2.8$ RQ and $+1.8$ mIoU without introducing additional computational burden.
Feature Interaction Methods. In Tab. 5, we explored the performance differences between Cross-Attention (CA) and Dynamic Depthwise Attention (DDA) two attention modules and their interaction with features of different scales. When using CA with multi-scale features only yields marginal improvements over our method, with a $+0.2$ increase in PQ and $+0.4$ in mIoU, but results in a significant slowdown, approximately 3.4×3.4$\times$3.4 × slower in FPS. Since single scale avoids the fuzzy spatial distribution caused by spatial dimension alignment, and the pre-perception of VAS can reduce the decoder interaction burden, using DDA with single-scale can bring a performance improvement of $+1.8$ PQ and $+2.4$ mIoU over DDA with multi-scale.
@@ Conclusion
We propose EOV-Seg, an efficient framework designed for open-vocabulary panoptic segmentation. It uses VAS and TDEE to improve instance discrimination capabilities for alleviating the interaction burden on mask decoder and endow mask recognition with spatial awareness capabilities. Extensive experiments show that our EOV-Seg is effective in terms of both accuracy and speed.
@@ Appendix A Overview
In the following supplementary materials, we elaborate on the inference details, provide the pseudocode of the Vocabulary-Aware Selection (VAS) module, and delineate the pipeline of the first two fusion methods of spatial awareness fusion, namely Early Aggregation Fusion (EAF) and Separable Dynamic Interaction (SDI). In addition, we present the additional experimental results and additional qualitative results, respectively.
@@ Appendix E Additional Experimental Results
Closed-Vocabulary and Open-Vocabulary panoptic segmentation performance. In Tab. 7, we compared EOV-Seg with SOTAs on COCO panoptic segmentation dataset to evaluate the closed-vocabulary capabilities. Specifically, with only a small increase in the number of parameters, the performance of EOV-Seg significantly surpassed FreeSeg (Qin et al. 2023), with PQ improved by $+15.5$, SQ improved by $+2.2$, RQ improved by $+18.4$, and mIoU improved by +18 and FPS is about 15×15$\times$15 × faster. Additionally, EOV-Seg outperforms ODISE (Xu et al. 2023a) by $+1.3$ PQ, $+7.8$ mIoU, and the inference speed is 17×17$\times$17 × faster with fewer parameters (-1297M). These results demonstrate that EOV-Seg not only has good generalization performance in open-vocabulary scenes, but also achieves excellent performance in closed-vocabulary scenes. In Tab. 8, we present open-vocabulary panoptic segmentation results on the Cityscapes (Cordts et al. 2016) dataset. EOV-Seg surpasses ODISE (Xu et al. 2023a), PQ is improved by $+10.8$, RQ is improved by $+13.9$ with fewer parameters (-1297M). Even under challenging street scene datasets, EOV-Seg has strong open-vocabulary recognition and segmentation capabilities.
@@
Method PQ SQ RQ mIoU FPS Params FreeSeg (Qin et al. 2023) 31.4 78.3 38.9 42.2 0.8 219M ODISE (Xu et al. 2023a) 45.6 - - 52.4 0.7 1522M EOV-Seg (L) 46.9 80.5 57.3 60.2 11.9 225M
@@
Method Scales PQ SQ RQ mIoU FPS EOV-Seg (S) (800, 1333) 14.3 54.1 17.7 20.4 16.4 (640, 2560) 15.1 57.0 18.9 21.9 24.5 EOV-Seg (L) (800, 1333) 23.1 70.2 28.3 31.3 9.1 (640, 2560) 24.5 70.2 30.1 32.1 11.6